OpenStudy (sesca89):

WILL FAN AND MEDAL!!! How do you find the area of a square? Isnt it like 2l x 2w??

jimthompson5910 (jim_thompson5910):

close, it's actually just the length times the width

OpenStudy (ajspeller):

\[side^2=Area\]

jimthompson5910 (jim_thompson5910):

So A = L*W A = area L = length W = width

jimthompson5910 (jim_thompson5910):

and as ajspeller is pointing out, you can say A = s^2 s = side length because all 4 sides are the same

OpenStudy (sesca89):

Well I have a problem and I'm ovethinking it, and I needed help for the formula. It's for polynomial functions and multiplying them to find a total area

OpenStudy (sesca89):

This is the problem

jimthompson5910 (jim_thompson5910):

what is the area of just the picture

OpenStudy (sesca89):

440

OpenStudy (sesca89):

but, i don't know the whole function

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):